OpenStudy (anonymous):

At noon, ship A is 170 km west of ship B. Ship A is sailing east at 40 km/h and ship B is sailing north at 15 km/h. How fast is the distance between the ships changing at 4:00 PM?

OpenStudy (anonymous):

What math is this?

OpenStudy (anonymous):

Calc 1

OpenStudy (anonymous):

Oh dang, no wonder. I tried for fun but I can't. Too high level for me. Sorry.

OpenStudy (anonymous):

its ok, thanks anyway!

OpenStudy (anonymous):

So after 4 hours ship A travels 4*40kms towards east (i.e. towards Ship B) = 160 Km So ship A is 10 Km West of Ship B now. Ship B travelled 15*4 Kms in 4 hours towards north so it travlled 60 Km North.

OpenStudy (anonymous):



OpenStudy (anonymous):

So by Pythagorean Theorem distance between them be x. x^2 = 60^2 + 10^2 so x = 60.83KM